# Generative Models From and For Sampling-Based MPC: A Bootstrapped Approach For Adaptive Contact-Rich Manipulation

### v1

@@ IV-C GPC-CEM: Bootstrapping SPC with Flow-Matching
      $\Sigma\leftarrow\mathrm{diag}\left(\mathrm{Var}(\{U^{(k)}\})\right)$
@@
Trained on a finite set of open-loop control sequences, the generative proposal distribution $p_{\theta}(U\mid\bm{x}_{\tau},\bm{h}_{\tau})$ inherits the generalization limitations of behavior cloning and model-based RL. This sensitivity to distributional shifts, something we also observe in our experiments with GPC-Shoot, manifests as degraded sample quality from the proposal distribution in a small subset of states, particularly when the object enters regions underrepresented in the training data. In contrast, SPC adapts its sampling distribution online and is more robust in such situations, though it remains myopic and computationally expensive. To balance these trade-offs, we bootstrap SPC with a flow-matching generative model that approximates the dataset’s control distribution while preserving the adaptability of online sampling during execution. We summarize our approach in Algorithm 1. At each control step, the algorithm begins by estimating the current state and shifting the current mean of the CEM sampling distribution forward in time. The key idea in GPC-CEM is to augment Gaussian CEM sampling with proposals from the generative model $p_{\theta}$ trained offline on control sequences that led to task success. Proposals yieling low-cost rollouts are used to update the CEM distribution $\pi_{\phi}(U)$, with the mean set to the time-shifted best candidate, and the variance is computed from the elite set, thereby adapting it toward higher-quality actions. Unlike standard CEM, the executed control is not derived from the elite mean, which assumes unimodality, but instead taken from the best-performing candidate, allowing GPC-CEM to better exploit multimodal proposals from the generative model. This hybrid approach efficiently guides exploration while maintaining the adaptability of online optimization.
@@ IV-D Application to Loco-Manipulation
We apply our bootstrapped SPC framework to non-prehensile object pushing with a Spot quadruped robot with an arm to demonstrate versatile loco-manipulation skills. With 19 degrees of freedom (DoF), planning for this robot is computationally expensive, especially in SPC, where higher dimensionality typically demands for larger sample sizes. To simplify the task, based on the work of [38], we disentangle the low-level locomotion control from the high-level task-level action space, as illustrated in Fig. 2. The high-level action space includes 9 DoF (3 for the torso, 6 for the arm) [^1] We exclude the gripper DoF from the high-level action space for non-prehensile manipulation. Yet, this and additional DoF (e.g., torso height, pitch, roll) could be included without retraining the low-level policy. and is mapped to low-level commands by a pre-trained locomotion policy that ensures balance and stability while tracking high-level inputs. The low-level locomotion policy is fixed throughout task planning and execution. Besides the lower-dimensional action space, a core advantage of this approach is that this hierarchical control structure naturally provides more robustness to the low-level control, removing the need to introduce guidance terms that explicitly enforce smoothness and temporal consistency of the control sequences into the flow-matching process.
Figure 2: The mapping from the high-level action space to the low-level spot loco-manipulation control space. The locomotion policy is pre-trained to follow the high-level commands while maintaining balance and stability. It is fixed throughout the full task planning and execution.
@@ V Experimental Results
We use conditional flow matching to train a generative model over SPC control sequences with a simple Multi-Layer Perceptron (MLP) as the underlying architecture. We consider both direct sampling from the learned model (GPC-Shoot) and the bootstrapped version that combines it with CEM-based online planning (GPC-CEM). We compare our approach against standard CEM, as well as model predictive path integral control (MPPI) [16] and DialMPC [18], a more recent SPC approach building on MPPI. We only compare against DialMPC on the Push-T task, as its inner optimization loop doubles simulation time, making it unsuitable for real-time loco-manipulation. All methods are implemented in Python using judo [39], which provides a unified interface for defining custom tasks and controllers. For each task, we evaluate all methods with the same number of rollouts per iteration, control frequency, and respective cost function. In addition, we report results for the GPC-methods across three different model seeds to account for the stochasticity in the training process. We set the CEM-sample ratio in GPC-CEM, i.e. $N_{\mathrm{CEM}}/N$, to 0.5 for both tasks.
@@
Fig. 4. Evaluation of intermediate models from the iterative training procedure from [14] on the Push-T task. Results are shown for 100 experiments per iteration.
@@ V-A Push-T Task
This task requires a 2-DoF circular robot to push a T-shaped block into a specified goal pose. Due to its sparse rewards and multi-modality, it serves as a popular benchmark for evaluating generative control policies. We also evaluate the adaptability of GPC to novel, unseen task variations without retraining by running it on a variant where the T-shaped block is replaced with a K-shaped block (Push-K), introducing different object dynamics. In this setting, we reuse the generative model trained on Push-T to bootstrap SPC for Push-K, showcasing GPC’s robustness and ability to generalize across task variations. Table I summarizes the simulation results. We report success rates with Wilson 95% confidence intervals and average completion times (for successful runs) with respective standard deviations based on 1000 trials per method. Success is defined as achieving at least 0.9 coverage of the target pose within 2500 time steps (0.01 s each). For GPC-CEM, we additionally report the average CEM sample ratio and standard deviation, indicating how often samples were selected from the CEM distribution over the learned proposal distribution.
Both CEM and DIAL-MPC [18] achieve high success rates ($\geq 85$%), demonstrating the strength of sampling-based methods. While DIAL-MPC includes an inner optimization loop, its gains over CEM are marginal and come at higher computational cost; thus, we use CEM for data collection. MPPI performs moderately (62% success), though better tuning may close this gap. GPC-Shoot performance improves with more denoising steps (indicated in parentheses): 99% success with 10 steps vs. 71% with 2. Similarly, GPC-CEM with 10 steps not only achieves 99.8% success but also reduces completion time by nearly 50% compared to 2 steps. Notably, GPC-CEM remains robust under reduced horizons (1s vs. 3s), maintaining 96% success versus 78% for CEM. This suggests that our method enables non-myopic planning under tighter computational constraints. In the Push-K generalization task, GPC-CEM again outperforms all baselines, achieving 96% success compared to 55% for CEM, indicating strong transferability of the learned proposal distribution. This is a notable insight, as CEM performance drops significantly without any changes to the cost function, highlighting that the learned distribution captures meaningful structural priors useful across tasks.
@@
We acknowledge the conceptual similarity between our approach and that of Kurtz et al. [14], who also combine generative modeling with SPC. However, their method adopts an iterative training procedure that alternates between data collection and model updates, similar to expert iteration in reinforcement learning [40]. This setup is motivated by the assumption that SPC data is too noisy to directly train a generative model; hence, each data collection iteration bootstraps SPC with a partially trained flow-matching model to improve the subsequent training distribution. In our experiments on the Push-T task, however, this iterative procedure did not improve performance (see Fig. I). In fact, success rates declined over training iterations. We attribute this to a reduction in multi-modality of the learned proposal distribution, which is critical for sampling-based MPC in tasks with multiple distinct solution modes. This interpretation is supported by the decreasing CEM sample ratio over training iterations, suggesting that the learned proposal increasingly matches the uni-modal CEM distribution, making the two distributions interchangeable rather than complementary. In contrast, our method trains a generative model directly on open-loop control sequences from SPC, without requiring iterative retraining. Despite the noisy data, our model achieves up to 99.2% success (GPC-Shoot with 10 denoising steps), and already reaches 96% when bootstrapped with CEM after a single training round.
@@ V-B Spot Loco-Manipulation
In this task, Spot must push a chair into a goal pose located behind a C-shaped obstacle (Fig. 5). The robot and chair are always initialized randomly on the opposite side of the obstacle, creating a local minimum that requires navigating around it to succeed. The task is further complicated by the high-dimensional action space and contact dynamics. Solving this with SPC demands long horizons and large sample sizes, both of which increase computational cost. A trial is considered successful if the chair’s position error is below 0.15 m and its yaw is within 50 degrees of the target.
@@
We first evaluate the task in simulation to enable larger-scale testing. Results are summarized in Table II. Each baseline is run for 100 trials, and GPC methods are evaluated with 3 model seeds (100 trials each). We omit DIAL-MPC due to its inner-loop optimization being too slow for real-time use in this task. As in Push-T, we report success rate, average completion steps (for successful runs), and CEM sample ratio. GPC-CEM outperforms all baselines, achieving up to 83% success with fewer executed steps. In contrast, CEM alone reaches only 33%, often failing due to limited horizon and sample budget. MPPI performs slightly better but remains unreliable under real-time constraints. GPC-CEM remains robust under reduced planning horizons (3 vs. 4 seconds), maintaining 60% success, while baseline performance degrades. This reinforces that learned proposals can enhance planning in resource-limited settings. Interestingly, fewer denoising steps (2 vs. 10) yield better performance in this task for both GPC-Shoot and GPC-CEM. We attribute this to reduced sample diversity at higher step counts, which impairs exploration in tasks with deceptive local minima. We also observe higher CEM sample ratios in this task compared to Push-T, indicating that the learned model alone (GPC-Shoot) is less accurate. Instead, it is most effective when used to augment CEM, highlighting the value of integrating learned proposals into online optimization rather than relying on them directly. Finally, we evaluate a task variant with an added obstacle avoidance cost to prevent collisions with the C-shaped obstacle. While omitted from the base task to avoid biasing the MPC methods, this constraint is essential for real-world deployment. In this setting, GPC-CEM still leads with a 56% success rate, outperforming MPPI (30%) and CEM (27%).
@@
We evaluate GPC-CEM and CEM on hardware including the obstacle avoidance cost in both cases. We rely on a Motion Capture system to track the current object state. GPC-CEM is run with 2 denoising steps for 20 trials, while CEM is limited to 10 trials due to frequent damaging failures to the robot, such as repeated collisions with the obstacle as it fails to navigate around. GPC-CEM achieves a 60% success rate (12/20), while CEM succeeds in only 10% of trials (1/10). Qualitative examples for GPC-CEM are shown in Fig. 5; all other runs are included in the supplementary video [^2] https://youtu.be/lKCGjjddv1E. CEM failures consistently result from getting stuck in the local minimum caused by the C-shaped obstacle, as it lacks the guidance from the generative model to sample motions that navigate around it. GPC-CEM only encounters this failure in 4 of 20 trials. The remaining failures stem from two causes: (1) pushing the chair beyond the workspace due to the lack of workspace constraints in the cost, and (2) discrepancies between simulated and real chair behavior, especially when assumptions about friction and contact are violated, such as when the chair’s wheels are unable to roll. We intentionally excluded workspace limits to evaluate the methods with minimal guidance.
@@
Finally, we analyze the computation time per MPC iteration, finding that the dominant bottleneck remains policy rollout, which accounts for over 90% of the total compute time, limiting the overall control frequency to 5 Hz. This overhead is primarily due to collision handling and contact dynamics in the physics engine.
VI Limitations
One limitation of our framework is that it does not explicitly address sim-to-real transfer, leaving it vulnerable to discrepancies between simulated and real-world dynamics. However, since our method relies on offline data collection, it can be trained on domain-randomized data, which we expect to improve robustness to variations in dynamics, actuator behavior, and sensor noise, as suggested by prior work [14]. In this work, all experiments consider fixed goals. We plan to extend our work to variable goals by transforming our data into goal-centric representations. The current system also does not use GPU acceleration for forward simulation or inference with the flow-matching model, but this could be addressed in future work to enable faster online planning and data collection, especially with larger sample sizes. Finally, the proposed approach is limited to state-based policies but can be extended to vision-based policies by learning from observations collected while executing the state-based policy in the real world or simulation. Future work should explore how to integrate vision-based action proposal distributions with fast vision-based dynamics models for online predictive control, similar to [13].
@@ VII-A Implementation Details
We describe the implementation of our generative model and sampling-based MPC algorithm, used for both offline data collection and online control. Code for data collection, model training and the bootstrapped SPC controller will be released upon acceptance.
@@
We collect training data using a sampling-based MPC controller in simulation for a maximum of 2500 time steps per episodes in both tasks, which generates open-loop control sequences. For Push-T, trajectories are cubic splines with 4 control points; for Spot, they consist of 4 linearly interpolated waypoints. All data is represented in the world frame, not relative to the robot. Although we tested robot-centric representations, they did not yield performance improvements. We collected 67,667 Push-T sequences from 1,000 successful episodes and 211,832 Spot sequences from 1,700 episodes. For evaluation, we use fixed sets of randomly sampled initial states for each task, shared across all methods and runs to ensure consistency.
@@
We train a conditional flow-matching model with simple MLPs, using a batch size of 40,000 and the Adam optimizer (learning rate 0.0001, cosine annealing schedule, 500 warmup steps) over 1,000 epochs. The model predicts 4 control points conditioned on the current robot and object state and the previous replanning state (history length = 1). Orientations are represented using sine-cosine encodings of yaw angles. Although Spot expects velocity commands, we predict absolute positions and convert them to velocities via finite differences during online control.
@@
Both tasks use a weighted sum of cost terms computed over the full-resolution control sequence (0.01s for Push-T, 0.02s for Spot). The cost components are:
  1. 1.
    Robot-Object Proximity: L2 distance between the robot and object. For Spot, we penalize both torso and end-effector distances.
  2. 2.
    Velocity Penalty: L2 norm of robot joint velocities.
  3. 3.
    Goal Reaching: L2 distance and angle difference between object and goal, plus a progress term penalizing lack of advancement over time.
  4. 4.
    Joint Limits (Spot): Penalty for exceeding arm joint limits (leg joints are handled by the low-level policy).
  5. 5.
    Fall Penalty (Spot): Large penalty if the torso height drops below a threshold, preventing falls.
  6. 6.
    Object Tipping (Spot): Large penalty if the chair’s z-axis deviates from vertical, discouraging it from tipping over.
Weights were tuned empirically and will be provided in the released code.

### v2

@@ IV-C GPC-CEM: Bootstrapping SPC with Flow-Matching
      $\Sigma\leftarrow\mathrm{diag}\left(\mathrm{Var}(\{U^{(k)}\}_{k=1}^{N_{\text{elite}}})\right)$
@@
Trained on a finite set of open-loop control sequences, the generative proposal distribution $p_{\theta}(U\mid\bm{x}_{\tau},\bm{h}_{\tau})$ inherits the generalization limitations of behavior cloning and model-based RL. This sensitivity to distributional shifts is something we also observe in our experiments with GPC-Shoot, where it manifests as degraded sample quality within regions underrepresented in the training data. In contrast, SPC adapts its sampling distribution online and becomes more robust in unseen situations, but remains myopic and computationally expensive. To balance these limitations, we bootstrap SPC with a flow-matching generative model that learns the dataset control distribution while preserving the adaptability of online sampling. We summarize our approach in Algorithm 1. At each control step, the algorithm begins by estimating the current state and shifting the current mean of the CEM sampling distribution forward in time. The key idea in GPC-CEM is to augment Gaussian CEM sampling with proposals from the generative model $p_{\theta}$ trained offline on control sequences that led to task success. The $N_{elite}$ proposals with the lowest-cost rollouts are used to update the CEM distribution’s $\pi_{\phi}(U)$ mean (the time-shifted lowest-cost proposal) and variance. Unlike standard CEM, the executed control is the best-performing candidate instead of the mean of the $N_{elite}$ proposals. This allows GPC-CEM to better exploit multimodal proposals from the generative model rather than collapsing them to a single modality, efficiently guiding exploration while maintaining the adaptability of online optimization.
@@ IV-D Application to Loco-Manipulation
We apply our bootstrapped SPC framework to non-prehensile object pushing with a Spot quadruped robot to demonstrate versatile loco-manipulation skills. With 19 positional degrees of freedom (DoF), planning for this robot is computationally expensive and typically demands large sample sizes. To simplify our sampling process, we disentangle low-level locomotion control based on the work of [42] and sample only in the high-level task action space. The high-level action space includes 9 DoF (3 for the torso, 6 for the arm) [^1] We exclude the gripper DoF from the high-level action space for non-prehensile manipulation, but this and additional DoFs (e.g., torso height, pitch, roll) could be included without retraining the low-level policy. and is mapped to the low-level commands by a pre-trained locomotion policy that ensures balance and stability while tracking high-level inputs. The low-level locomotion policy is fixed throughout task planning and execution. In addition to a lower-dimensional action space, this hierarchical control structure naturally provides more robustness to the low-level control. As a result, we do not need to explicitly enforce strong smoothness or temporal consistency constraints in the high-level action space used for flow matching.
@@ V Experimental Results
We use conditional flow matching to train a generative model over SPC control sequences with a Multi-Layer Perceptron (MLP) as the underlying architecture. We also baseline the flow model against a CVAE trained on the same data. We consider both direct sampling from the learned model (GPC-Shoot) and the bootstrapped version that combines it with CEM-based online planning (GPC-CEM). We compare our approach to standard CEM, model predictive path integral control (MPPI); [38]; and DialMPC [39], a more recent SPC approach building on MPPI. We only compare against DialMPC and GPC-Shoot with a CVAE on Push-T; the former’s inner optimization loop makes it unsuitable for our real-time loco-manipulation examples, while the latter proved inferior to flow matching. All methods are implemented in Python using judo [23] as a unified interface for defining custom tasks and controllers. For each task, we evaluate all methods with the same number of rollouts per iteration, control frequency, and respective cost function. In addition, we report results for the GPC-methods across three different model seeds to account for the stochasticity during training. We set the CEM-sample ratio in GPC-CEM, i.e. $N_{\mathrm{CEM}}/N$, to 0.5 for both tasks.
@@ V-A Push-T Task
This task requires a 2-DoF circular robot to push a T-shaped block to a specified goal pose. Due to its sparse rewards and multi-modality, it serves as a popular benchmark for evaluating generative control policies. We also evaluate the adaptability of GPC to unseen task variations by running it on a variant using a K-shaped block (Push-K) with different object dynamics. In this setting, we reuse the generative model trained on Push-T to bootstrap SPC for Push-K without retraining, showcasing GPC’s ability to generalize across task variations. Table I summarizes the simulation results. We report success rates with Wilson 95% confidence intervals and average completion times (for successful runs) with respective standard deviations based on 1000 trials per method. Success is defined as achieving at least 90% coverage of the targetxx pose within 2500 time steps (0.01 s each). For GPC-CEM, we additionally report the average CEM sample ratio and standard deviation, indicating how often samples were selected from the CEM distribution over the learned proposal distribution.
Both CEM and DIAL-MPC [39] achieve high success rates ($\geq 85$%), demonstrating the strength of sampling-based methods. DIAL-MPC’s gains over CEM are marginal and come at higher computational cost, so we use CEM for data collection. MPPI achieves 62% success, though better tuning may close this gap. Flow-based GPC-Shoot improves with more denoising steps (indicated in parentheses): 99% success with 10 steps vs. 71% with 2, while CVAE-based GPC-Shoot achieves 97% success. Similarly, GPC-CEM with 10 steps not only achieves 99.8% success but also reduces completion time by nearly 50% compared to 2 steps. Notably, GPC-CEM remains robust under reduced horizons (1s vs. 3s), maintaining 96% success versus 78% for CEM. This suggests that our method enables non-myopic planning under tighter computational constraints. In the Push-K generalization task, GPC-CEM again outperforms all baselines, achieving 96% success compared to 55% for CEM, indicating strong transferability of the learned proposal distribution. This is a notable insight, as both CVAE-based GPC-Shoot and CEM performance drops significantly without any changes to the cost function, highlighting that the flow-based learned distribution captures structural priors useful across tasks.
Figure 3: Push-T Simulation Studies. Left: Evaluation of intermediate models from the iterative training procedure of [22]. Right: Ablation study on sample split.
@@
We acknowledge the conceptual similarity between our approach and that of Kurtz et al. [22], who also combine generative modeling with SPC. However, their method adopts an iterative training procedure that alternates between data collection and model updates, similar to expert iteration in reinforcement learning [1]. This setup is motivated by the assumption that SPC data is too noisy to directly train a generative model; hence, each data collection iteration bootstraps SPC with a partially trained flow-matching model to improve the subsequent training distribution. In our experiments on the Push-T task, however, this iterative procedure did not improve performance (see Fig. 3). In fact, success rates decline over training iterations. In a qualitative analysis, we find that the resulting policies tend to collapse to small, random movements that fail to complete the task. We interpret this as the iterative training procedure gradually diminishing the multi-modality of the learned proposal distribution. Consistent with this interpretation, we also observe a decreasing CEM sample ratio over training iterations, suggesting that the learned proposal distribution converges to mimic the CEM sampling distribution and reduces their complementarity. In contrast, our method trains a generative model directly on open-loop control sequences from SPC, without requiring iterative retraining. Despite the noisy data, our model achieves up to 99.2% success (GPC-Shoot with 10 denoising steps) and already reaches 96% when bootstrapped with CEM after a single training round.
Ablations
We conduct three ablation studies to further analyze the performance of GPC-CEM and GPC-Shoot on the simulated Push-T task. Ablation 1 measures the effect of the sample split between samples drawn from the CEM distribution and the learned proposal distribution. We vary the sample split $N_{\mathrm{CEM}}/N$ from 0.1 to 1.0 (only CEM). As shown in Fig. 3, we find that increasing the number of CEM samples improves or maintains the success rate until exclusively using CEM samples, where performance drops significantly. This highlights that the learned proposal distribution contributes valuable samples that both complement and strengthen the CEM distribution. We further observe that the empirical CEM sample ratio grows sublinearly with respect to the specified split, indicating that even a relatively small fraction of learned proposal samples disproportionately contribute to the overall sampling process. Ablation 2 analyzes the impact of the dataset size used to train the flow-matching model. To assess how many demonstrations are needed to train an effective proposal model, we vary the dataset size from 100 to 1000 MPC rollouts [^2] Each rollout corresponds to a single trajectory of $H/\Delta t$ steps.. Performance improves rapidly with data and surpasses 90% success with only 200 rollouts, after which performance gains steadily saturate. This shows that the proposal model can be trained in a sample efficient manner requiring just a few hundred trajectories, whereas reinforcement learning methods typically need orders of magnitude more interaction to achieve comparable performance. Ablation 3 assesses model architecture by comparing flow model performance against a conditional variational auto-encoder (CVAE) using GPC-Shoot. Table I shows the ten-step flow model outperformed the CVAE for each task and variation with a similar number of steps. Along with its degraded performance on the Push-K task, the CVAE’s struggle to adjust to a different horizon length was characteristic of its documented difficulties with multi-modal generalization and domain adaptation [9]. Due to its higher performance and robust generalization, we selected flow matching as our generative model.
@@ V-B Spot Loco-Manipulation
In this task, Spot must push a chair to a goal pose located behind a C-shaped obstacle (Fig. 4). The robot and chair are always initialized randomly on the opposite side of the obstacle, creating a local minimum that requires navigating around it to succeed. The task is further complicated by the high-dimensional action space and contact dynamics. Solving this with SPC requires long horizons and large sample sizes, both of which increase computational cost. A trial is considered successful if the chair’s position error is below 0.15 m and its yaw is within 50 degrees of the target.
@@
We first evaluate the task in simulation to enable larger-scale testing. Results are summarized in Table II. Each baseline is run for 100 trials, and GPC methods are evaluated with 3 model seeds (100 trials each). We omit DIAL-MPC due to its inner-loop optimization being too slow for real-time use in this task. As in Push-T, we report success rate, average completion steps (for successful runs), and CEM sample ratio. GPC-CEM outperforms all baselines, achieving up to 83% success with fewer executed steps. In contrast, CEM alone reaches only 33%, often failing due to limited horizon and sample budget. MPPI performs slightly better but remains unreliable under real-time constraints. GPC-CEM remains robust under reduced planning horizons (3 vs. 4 seconds), maintaining 60% success while baseline performance degrades. This reinforces that learned proposals can enhance planning in resource-limited settings. Interestingly, fewer denoising steps (2 vs. 10) yield better performance in this task for both GPC-Shoot and GPC-CEM. We attribute this to reduced sample diversity at higher step counts, which impairs exploration in tasks with deceptive local minima. We also observe higher CEM sample ratios in this task compared to Push-T, indicating that the learned model alone (GPC-Shoot) is less accurate. Instead, it is most effective when used to augment CEM, highlighting the value of integrating learned proposals into online optimization rather than relying on them directly. Finally, we evaluate a task variant with an added obstacle avoidance cost to prevent collisions with the C-shaped obstacle. This is omitted from the base task to avoid biasing the MPC methods, but is essential for real-world deployment. In this setting, GPC-CEM still leads with a 56% success rate, outperforming MPPI (30%) and CEM (27%).
@@
We evaluate GPC-CEM and CEM on hardware including the obstacle avoidance cost in both cases. We rely on a Motion Capture system to track the object state. GPC-CEM is run with 2 denoising steps for 20 trials, while CEM is limited to 10 trials due to frequent damaging failures to the robot (e.g. repeated collisions with the obstacle as it fails to navigate around). GPC-CEM achieves a 60% success rate (12/20), while CEM succeeds in only 10% of trials (1/10). Qualitative examples for GPC-CEM are shown in Fig. 4; all other runs are included in the supplementary video [^3] https://youtu.be/lKCGjjddv1E. CEM failures consistently result in the local minimum caused by the C-shaped obstacle, as it lacks the guidance from the generative model to sample motions that navigate around it. GPC-CEM only encounters this failure in 4 of 20 trials. The remaining failures stem from two causes: (1) pushing the chair beyond the workspace due to the lack of workspace constraints in the cost, and (2) discrepancies between simulated and real chair behavior, especially assumptions about friction and contact such as when the chair’s wheels can roll.
@@
We find that the policy rollout accounts for over 90% of the total compute time and becomes the primary computation bottleneck, limiting the overall control frequency to 5 Hz. This overhead is primarily due to collision handling and contact dynamics in the physics engine.
VI Limitations and Future Work
Our framework does not explicitly address sim-to-real transfer, leaving it vulnerable to discrepancies between simulated and real-world dynamics. However, since it relies on offline data collection, it can be trained on domain-randomized data to improve robustness to variations in dynamics, actuator behavior, and sensor noise [22]. In addition, learning proposal distributions from a combination of simulated and real-world data could further enhance transferability and performance during hardware deployment. In this work, all experiments consider fixed goals in a world frame. We plan to extend our work to variable goals by transforming our data into goal-centric representations. The current system also does not use GPU acceleration for simulation or proposal inference, but this could be addressed in future work to enable faster online planning and data collection (particularly with larger sample sizes). Finally, the proposed approach is limited to state-based policies but can be distilled to vision-based policies by learning from observations collected while executing the state-based policy in the real world or simulation. Future work can explore how to integrate vision-based action proposal distributions with fast vision-based dynamics models for online predictive control [30]. Although our offline data collection already captures long-horizon structure by using extended SPC horizons, future work could also incorporate learned infinite-horizon value functions. Such value estimates would provide an additional source of global, task-level guidance, while our generative priors would continue to shape and improve the search over control sequences during online optimization.
@@ VII-A Implementation Details
We describe the implementation of our generative model and sampling-based MPC algorithm, used for both offline data collection and online control.
@@
We collect training data using a sampling-based MPC controller in simulation for a maximum of 2500 time steps per episodes in both tasks, which generates open-loop control sequences. For Push-T, trajectories are cubic splines with 4 control points; for Spot, they consist of 4 linearly interpolated waypoints. All data is represented in the world frame, not relative to the robot. Although we tested robot-centric representations, they did not yield performance improvements. For Spot, the manipulated object is modeled as a single free rigid body with empirically tuned mass, inertia, and friction, using simplified collision geometry and low base friction to approximate rolling behavior. We collected 67,667 Push-T sequences from 1,000 successful episodes and 211,832 Spot sequences from 1,700 episodes. For evaluation, we use fixed sets of randomly sampled initial states for each task, shared across all methods and runs to ensure consistency.
@@
We implemented our baseline CVAE and conditional flow-matching models as MLPs trained with a batch size of 40,000 and the Adam optimizer with a learning rate 0.0001, cosine annealing schedule, 500 warmup steps over 1,000 epochs. The model predicts 4 control points conditioned on the current robot and object state and the previous replanning state (history length = 1). Orientations are represented using sine-cosine encodings of yaw angles. Although Spot expects velocity commands, we predict absolute positions and convert them to velocities via finite differences during online control. Weights were tuned empirically.
Training Details
We implemented our baseline CVAE and conditional flow-matching models as MLPs trained with a batch size of 40,000 and the Adam optimizer with a learning rate 0.0001, cosine annealing schedule, 500 warmup steps over 1,000 epochs. The model predicts 4 control points conditioned on the current robot and object state and the previous replanning state (history length = 1). Orientations are represented using sine-cosine encodings of yaw angles. Although Spot expects velocity commands, we predict absolute positions and convert them to velocities via finite differences during online control. Weights were tuned empirically.
@@
Both tasks use a weighted sum of costs computed over the full-resolution control sequence (0.01s for Push-T, 0.02s for Spot). The cost components include robot–object proximity (L2 distance between the robot and object, penalizing both torso and end-effector distances for Spot), a velocity penalty given by the L2 norm of robot joint velocities, goal reaching terms measuring L2 distance and angle difference between the object and goal with an additional progress penalty, joint limit penalties for exceeding arm joint limits on Spot (leg joints handled by the low-level policy), a fall penalty applied when the Spot torso height drops below a threshold, and an object tipping penalty when the chair’s z-axis deviates from vertical.